dad is seven times as old as his son. 10 years later, he will be three times as old as his son. find their ages.
Answer by mananth(16946)

let son be x years
dad = 7x
..
10 years later
son = x+10 years
dad = 7x +10 years
..
7x+10=3(x+10)
7x+10=3x+30
7x-3x=30-10
4x = 20
x= 5
son's age now = 5 years
dad's age now = 7*5 = 35 years
..